A process is in-control and centered at 440 with σ = 1.5. The process is normally distributed with a SQL = 5.25. What are the upper and lower specifications limits for the process?
Answer #1

Specification limit= centre value ± sigma* SQL

SO, Upper specification limit=440+1.5*5.25=447.875

Lower specification limit=440-1.5*5.25= 432.125